# Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism

### v1

@@ 3.2. Offline Profiler
3.3. Transformer IR
APEX identifies an optimal parallel execution plan by generating and evaluating multiple candidate execution plans. Yet, it is non-trivial to generate parallel execution plans for a variety of LLMs as they have different model architectures. To address this, we introduce the Transformer IR, a unified abstraction for transformer-based models. APEX utilizes Parallel Templates, developed based on the Transformer IR rather than specific models, to generate parallel execution plans. This approach enables APEX to parallelize a wide range of LLMs that can be represented through the Transformer IR. We discuss the details of Parallel Templates in Section 3.4.
We depict the idea of Transformer IR in Figure 3. We define the key operations of transformers, such as the multi-head attention, as cells. An LLM can be represented as a chain of cells. For example, GPT-3 (Brown et al., 2020) models can be represented as a chain of multi-head attention cells and multi-layer perceptron cells, and Llama-3.1 (Dubey et al., 2024) models can be represented as a chain of group query attention (Ainslie et al., 2023) cells and SwiGLU (Shazeer, 2020) cells. The Transformer IR represents an LLM canonically, which only captures the key cells in the model and ignores operations like tokenization and position embeddings, as they are less relevant for model parallelization; this reduces the search space for simulation. Given that LLMs often use the same set of cells repeatedly, we further define the smallest set of non-repetitive adjacent cells as a block. Each cell consists of multiple tasks, such as the attention heads in the multi-head attention cell, or the experts in the mixture-of-expert cell. Each task works independently from the other tasks (i.e., no inter-communication is needed), and their outputs are combined via operations like concatenation or AllReduce. To summarize, using the Transformer IR, an LLM can be unifiedly represented as multiple identical blocks, each block consists of numerous cells, and each cell consists of multiple tasks. We develop an IR converter, which parses the information of an LLM configuration file, such as the number of attention heads, number of layers, etc., and represents the model using Transformer IR.
3.4. Parallel Templates
We develop Parallel Templates to generate various parallel execution plans for LLM serving. As mentioned in Section 3.3, the templates are developed based on the Transformer IR, which allows the templates to support a wide range of LLMs. Each cell type is associated with a pre-defined Parallel Template that describes how it can be parallelized on multiple devices. We show some examples in Figure 4. If tensor parallelism is applied to both the multi-head attention (MHA) cell and the multi-layer perceptron (MLP) cell, as in Figure 4 (a), the tasks (e.g., attention heads) in the two cells are evenly distributed to each device. In addition to distributing the tasks, the Parallel Templates also handle the collective communications between adjacent cells. In the example of Figure 4 (a), an AllReduce is performed between the cells for data synchronization, which is similar to the case in Megatron-LM (Shoeybi et al., 2020). If the numbers of cell replicas vary between the adjacent cells, tensor resharding is required for the output activation; such cases happen when different data parallelism (DP) degrees are applied. For example, in Figure 4 (b), a 2-way DP is applied to the MLP cell, and a 1-way DP (i.e., no DP) is applied to the MHA cell, resulting in a different number of cells replicas. Consequently, All-to-All and AllGather operations are performed for tensor resharding. Figure 4 (c) depicts a Parallel Template associated with the MHA cell: $H$ attention heads are evenly distributed to $D$ device for computation, and collective communications for resharding and synchronization may also be performed, depending on the number of cell replicas of the adjacent cells. Note that while Figure 4 shows a simplified example of having only two devices, the Parallel Templates are parameterized and can support any number of devices.
@@ 3.4. Parallel Templates
3.5. Parallel Scheme Generator & Device Mapper
Given an LLM and pre-defined Parallel Templates, the Parallel Scheme Generator produces various parallel schemes, which maps the LLM onto a logical device cluster (Zhao et al., 2020) with various combinations of parallelism techniques. A logical device cluster is a virtual overlay that describes the number of devices in the cluster without the information of the network topology. Consequently, a parallel scheme only defines the parallelisms applied to the Transformer IR cells and their corresponding collective communications, such as the examples in Figure 4 (a) and (b), without assigning the cells to specific physical devices. We introduce a virtual overlay of a logical device cluster for design simplicity; this approach allows us to decouple the Parallel Template designs (Section 3.4) from the hierarchal network topology of a device cluster - jointly considering the network topology in the Parallel Templates would greatly increase design complexity. The Parallel Scheme Generator creates parallel schemes using a hierarchical top-down approach, beginning with the most coarse-grained model-level parallelism and gradually refining to fine-grained cell-level parallelism. The detailed procedure is outlined in Algorithm 1. First, the Generator decides the degree of model-level data parallelism (i.e., the number of model replicas), and also derives the number of devices assigned to each replica (i.e., number of replica devices). APEX requires the device cluster to be evenly partitioned by the number of model replicas, so the available degrees of parallelism are restricted to the divisors of the total number of devices in the cluster. Second, the Generator decides the number of pipeline stages within each replica. Similarly, APEX requires the cluster to be evenly partitioned by the number of stages, so the available parallelism degrees are the divisors of the number of devices assigned to the replica. Third, the Generator decides the degree of cell-level data parallelism for each cell, and generates task mappings using the pre-defined Parallel Templates. If the degree of cell-level data parallelism for a specific cell is less than the number of assigned devices, the Parallel Templates apply intra-layer parallelism like tensor or expert parallelism to parallelize the cell (Section 3.4). For example, if four devices are assigned to a cell with two replicas, two devices are assigned to each replica, and the cell can be parallelized using tensor parallelism. Afterward, for each pair of adjacent cells, the Generator inserts collective communications into the parallel schemes if tensor resharding is needed.
@@ 3.5. Parallel Scheme Generator & Device Mapper
After generating the parallel schemes, the Device Mapper maps the logical devices onto the physical device cluster, producing parallel execution plans. While the parallel schemes are generated in a top-down approach, the Device Mapper operates in a bottom-up manner. The logical devices are first mapped to physical devices connected at the bottom level of the cluster (see Figure 1); if the number of logical devices exceeds the available physical devices at the current level, the Device Mapper moves to the next upper level to include additional physical devices for mapping. Since lower-level device connections generally offer higher bandwidth than upper-level connections, the Device Mapper prioritizes mapping logical devices assigned to the same cell, as fine-grained cell-level parallelisms tend to incur expensive collective communication overhead, such as the AllReduce in tensor parallelism. The Device Mapper then maps logical devices assigned to the same pipeline stage, followed by those assigned to the same model replica, progressing through increasingly coarse-grained parallelism levels. This bottom-up approach maps logical devices with potentially higher communication overhead to physical devices connected at the lower level, which have higher interconnection bandwidth, and vice versa.
3.6. Iteration-Level Batching Module
@@
The Iteration-Level Batching Module takes the candidate execution plans and request traces as inputs, and simulates the behavior of a serving system using iteration-level batching (Yu et al., 2022). Each request comprises three key attributes: context length, generation length, and arrival time. As mentioned in Section 2.2, iteration-level batching continuously adds arriving requests to the current batch whenever memory permits. Therefore, a Resource Monitor keeps track of the memory usage and determines whether a new request can be batched. The Request Scheduler also checks the arrival time of a request to determine whether a request has arrived and can be batched. The Request Scheduler maintains a list of active requests, which are the requests in the current batch. The active request list also tracks how many tokens have been generated for each request. If the number of generated tokens matches the generation length of a request, then the request is completed and should be removed from the active batch; this releases memory to batch new requests. One token is generated for each active request in one iteration. In each iteration, the Iteration-Level Batching Module sends the active request list to the Dynamism-Aware Simulator (Section 3.7) to estimate the execution time. The Request Scheduler greedily adds arriving requests to the batch whenever memory is available to accommodate the context length of a request, without considering the memory required for the KV cache of generated tokens. As a result, memory capacity may be exhausted before token generation for all requests is completed. In such cases, the most recently added requests and their generated tokens are temporarily removed from the batch to free up memory for earlier requests to complete their token generation. These removed requests are re-added to the batch as memory becomes available. Both the removal and re-addition are performed in order, prioritizing completing the earliest requests.
3.7. Dynamism-Aware Simulator
Given the active requests produced by the Iteration-Level Batching Module and the Profiling Results, the Dynamism-Aware Simulator estimates the execution time of each iteration. The active request list contains information such as the context length of the request and the number of tokens that have already been generated for a request. This allows the Simulator to identify whether a request is in the prefill stage or the generation stage. Specifically, a request is in the prefill stage if the number of generated tokens equals zero; otherwise, it is in the generation stage. For requests in the prefill stage, we calculate the execution time of the operations by setting the input sequence length as the context length, as the entire sequence is processed in parallel (Section 2.1). For requests in the generation stage, we calculate the execution time of the operations by setting the input sequence length as one; this is because the output states (i.e., key and value) of the previous tokens are stored in the KV cache and do not require recomputation. Thus, the computation is only performed on the current token, reducing the effective sequence length to one. Furthermore, requests in the generation stage are processed in parallel; thus, assuming $n$ generation requests, the Simulator estimates the execution time by setting the sequence length as $n$, as each request has an effective sequence length of one. The Profiling Results contain the execution time of different attention heads, hidden dimensions, sequence lengths, etc., allowing the Simulator to estimate the total execution time for processing the prefill requests and the generation requests. If a specific data point is not presented in the Profiling Results, the Simulator estimates its value by leveraging linear interpolation between the nearest available profiling points. The Simulator is dynamism-aware as the Iteration-Level Batching Module continuously updates the list of active requests in every iteration. In addition, the Simulator utilizes the repetitive nature of LLMs to reduce simulation overhead. As discussed in Section 3.3, a block in Transformer IR is the smallest set of non-repetitive cells. Thus, the Simulator performs simulation using a single block and projects the execution time of the entire model, which consists of multiple repetitive blocks. The projection calculation, such as summing or taking the maximum of the execution times, depends on the connection pattern of the blocks (e.g., sequential or pipelined). In addition to estimating the execution time of each iteration, the Simulator also features a Metric Monitor, which calculates the value of several important metrics in LLM serving systems, such as time to first token (TTFT), time per output token (TPOT), P95 latency, Model FLOPs Utilization (MFU), Model Bandwidth Utilization (MBU), among others.
@@ 4. Experimental Results
4.1. Experimental Setup
To evaluate APEX, we choose a broad range of LLMs and datasets with distinct workloads. We also experiment with different data types. We discuss the setup details below.
Models: For evaluation, we choose four state-of-the-art LLMs: Qwen2.5-32B-Instruct (Hui et al., 2024), Llama-3.1-70B-Instruct (Dubey et al., 2024), Mistral-Large-Instruct (123B) (Jiang et al., 2023), and Mixtral 8x22B (Jiang et al., 2024); This covers a myriad of LLMs in different sizes, ranging from 32B to 123B, and also a Mixture-of-Expert (MoE) model. While the models use half-precision (FP16) by default, we also test cases where the model is quantized to FP8 format.
Datasets: We prepared three sets of request traces derived from distinct datasets: a paper abstract summarization dataset (Cohan et al., 2018), a news abstract summarization dataset (Fabbri et al., 2019), and a conversational dataset, LMSYS-Chat-1M (Zheng et al., 2024). These datasets represent three distinct workloads. The paper abstract summarization (Cohan et al., 2018) exemplifies a prefill-intensive workload, which has a long context length and short generation length. While some LLM workloads, such as paper summarization, are prefill-intensive, some (e.g., code generation and storytelling) are generation-intensive, involving short context lengths and long generation lengths. To create a generation-intensive workload, we adapt the news abstract summarization dataset (Fabbri et al., 2019), which consists of short summarizations. We modified the dataset by appending the following prompt to each news summarization: “Please generate a long story using the provided abstract,” and use them as input. This prompts the LLM to produce a long story from a short abstract, resulting in a generation-intensive workload. The LMSYS-Chat-1M (Zheng et al., 2024) dataset collects a series of real-world conversations between users and LLMs. Most requests in this dataset feature short context and generation lengths, representing a lightweight conversational workload. For evaluation, we randomly subsampled between 512 and 1K requests from each dataset. To simplify terminology, we refer to these datasets as Summarization (paper abstracts), Creation (news generation), and Chat (LMSYS-Chat-1M) in the following sections. We assume a Poisson distribution for request arrival times. Details of the request traces are listed in Table 1. We report context lengths and generation lengths in terms of tokens, which vary for different tokenizers. We use the Llama-3.1 tokenizer as an example in Table 1; other tokenizers produce similar token counts.
Hardware Platform and Serving Systems: We run APEX simulation using an Intel Xeon 6530 CPU and validated the simulation results against the actual performance of LLM serving on a GPU cluster. We deployed vLLM (Kwon et al., 2023) v0.6.0 to configure an LLM online server on a GPU cluster equipped with 8 Nvidia H100 SXM GPUs, each with 80 GB of GPU memory. Requests were sent to the server according to their arrival times. Since vLLM does not natively support data parallelism, we implemented data parallelism by setting up multiple vLLM servers and dispatching requests to them in a round-robin fashion.
@@ 4.1. Experimental Setup
Summarization Qwen-2.5-32B 0.25 2153.27 (1$\times$) 1137.70 (1.89$\times$) 968.26 (2.22$\times$)
0.5 1823.85 (1$\times$) 1116.84 (1.63$\times$) 967.23 (1.89$\times$)
Llama-3.1-70B 0.25 1998.82 (1$\times$) 1340.99 (1.49$\times$) 1175.09 (1.70$\times$)
@@
Mistral-Large 0.25 1301.27 (1$\times$) 498.32 (2.61$\times$) 376.13 (3.46$\times$)
0.5 1140.36 (1$\times$) 498.32 (2.29$\times$) 376.13 (3.03$\times$)
Mixtral-8x22B 0.25 1841.90 (1$\times$) 1290.97 (1.43$\times$) 551.94 (3.34$\times$)
0.5 1429.46 (1$\times$) 1137.87 (1.26$\times$) 552.28 (2.59$\times$)
Creation Qwen-2.5-32B 0.25 5718.22 (1$\times$) 2432.1 (2.35$\times$) 1507.68 (3.79$\times$)
0.5 4950.96 (1$\times$) 2421.82 (2.04$\times$) 1512.95 (3.27$\times$)
Llama-3.1-70B 0.25 6549.85 (1$\times$) 3742.77 (1.75$\times$) 3159.96 (2.07$\times$)
0.5 5240.78 (1$\times$) 4578.49 (1.14$\times$) 3183.67 (1.65$\times$)
Mistral-Large 0.25 6379.09 (1$\times$) 4063.11 (1.57$\times$) 2684.37 (2.38$\times$)
0.5 5221.52 (1$\times$) 4121.53 (1.27$\times$) 2682.05 (1.95$\times$)
Mixtral-8x22B 0.25 2645.24 (1$\times$) 2423.82 (1.09$\times$) 1039.87 (2.54$\times$)
0.5 1442.15 (1$\times$) 1417.75 (1.02$\times$) 1035.62 (1.39$\times$)
Chat Qwen-2.5-32B 0.25 1251.75 (1$\times$) 798.75 (1.57$\times$) 513.38 (2.44$\times$)
0.5 1115.01 (1$\times$) 786.08 (1.42$\times$) 510.72 (2.18$\times$)
Llama-3.1-70B 0.25 1622.48 (1$\times$) 1118.85 (1.45$\times$) 824.43 (1.97$\times$)
@@
Mistral-Large 0.25 973.16 (1$\times$) 396.32 (2.46$\times$) 254.26 (3.83$\times$)
0.5 890.71 (1$\times$) 391.96 (2.27$\times$) 254.26 (3.50$\times$)
Mixtral-8x22B 0.25 1757.68 (1$\times$) 1504.88 (1.17$\times$) 397.93 (4.42$\times$)
0.5 1406.75 (1$\times$) 1253.22 (1.12$\times$) 396.80 (3.55$\times$)
4.2. Evaluation of APEX’s suggestion

### v2

@@ 3.2.1. IR Converter and Transformer IR
3.2.2. Parallel Templates and Parallel Schemes Generator
We develop Parallel Templates to generate parallel execution plans for LLM serving. Given the wide variety of LLMs, exhaustively designing templates for each model is impractical. Instead, we develop our Parallel Templates based on the Transformer IR, which provides a unified abstraction of LLMs. This approach enables the developed templates to support a broad range of models rather than being tailored to specific ones. Figure 5 illustrates examples of our templates. Each cell type is associated with a pre-defined template that specifies how it is parallelized across multiple devices. For instance, if tensor parallelism (TP) is applied to both the multi-head attention (MHA) cell and the multi-layer perceptron (MLP) cell, as shown in Figure 5 (a), tasks (e.g., attention heads) are evenly distributed across devices. Beyond task distribution, Parallel Templates also manage collective communication between adjacent cells. In Figure 5 (a), an AllReduce operation synchronizes data between cells, similar to the approach in Megatron-LM (Shoeybi et al., 2020). In Figure 5 (b), 2-way data parallelism (DP) is applied to the MLP cell, while no DP is applied to the MHA cell, leading to distinct model partitioning. Consequently, All-to-All and AllGather operations are performed for tensor resharding. While Figures 5 (a) and (b) depict simplified examples using two devices, Parallel Templates are fully parameterized to support any number of devices, as in Figure 5 (c), where $H$ attention heads are evenly distributed across $D$ devices.
@@ 3.2.2. Parallel Templates and Parallel Schemes Generator
Given Parallel Templates and an LLM represented in Transformer IR, the Parallel Scheme Generator constructs parallel schemes that map the LLM onto a logical device cluster (Zhao et al., 2020) using different combinations of parallelism techniques. A logical device cluster is a virtual overlay that specifies the number of devices in a cluster without considering network topology. As a result, a parallel scheme defines the parallelism applied to Transformer IR cells and their collective communications (e.g., Figure 5 (a), (b)) without assigning the cells to specific physical devices. Introducing logical device clusters simplifies design complexity by decoupling Parallel Templates from the hierarchical network topology of physical clusters; the Device Mapper later handles physical mappings. The Parallel Scheme Generator follows a hierarchical top-down approach, starting with coarse-grained model-level parallelism and progressively refining to fine-grained cell-level parallelism. As detailed in Algorithm 1, the Generator first selects the degree of model-level data parallelism (i.e., number of model replicas) and partitions the cluster accordingly. APEX enforces even partitioning, restricting parallelism choices to divisors of the total number of devices. Next, the Generator determines the number of pipeline stages within each replica, again ensuring even partitioning. Finally, it assigns cell-level data parallelism and applies Parallel Templates to generate task mappings. If a cell’s data parallelism degree is lower than the number of assigned devices, intra-layer parallelism (e.g., tensor or expert parallelism) is applied. For instance, if four devices are assigned to a cell with two replicas, each replica receives two devices, and tensor parallelism is applied to parallelize the cell replica across the two devices. Lastly, for each pair of adjacent cells, the Generator inserts collective communication operations for synchronization and tensor resharding. Now that various parallel schemes have been generated for the target model, the logical devices can be mapped to physical devices, producing executable parallel execution plans; we discuss the physical device mapping in the following subsection.
@@ 3.2.3. Device Mapper
The Device Mapper takes the parallel schemes and the device cluster configuration as input, and assigns logical devices cluster to the physical device cluster, producing parallel execution plans. While the parallel schemes are generated in a top-down approach, the Device Mapper operates in a bottom-up manner. Specifically, the logical devices are first mapped to physical devices connected at the bottom level of the cluster, such as Level 1 in Figure 1; if the number of logical devices exceeds the available physical devices at the current level, the mapping extends to the next higher level to include additional physical devices. Since lower-level connections generally offer higher bandwidth (e.g., NVLink in Level 1), the Device Mapper prioritizes mapping logical devices assigned to the same cell, as fine-grained parallelism (e.g., tensor parallelism) often incurs expensive collective communication (e.g., AllReduce). Next, the Device Mapper maps logical devices assigned to the same pipeline stage, followed by those assigned to the same model replica, progressing through increasingly coarse-grained parallelisms. This bottom-up approach ensures that logical devices with higher communication demands are mapped to lower-level physical connections with greater bandwidth, while those with lower communication overhead are placed at higher levels. Once the mapping is completed, we obtain multiple parallel execution plans, each describing a distinct strategy for parallelizing the input LLM on the target cluster. Next, these plans are evaluated using the Batching Module and the LLM Serving Simulator.
3.3. Batching Module
@@
The Batching Module takes the parallel execution plans and a trace of requests as inputs, and simulates LLM serving system using iteration-level batching (Yu et al., 2022). Request traces are often collected from online services like Azure LLM Inference traces (Patel et al., 2024), and each request comprises three key attributes: context length, generation length, and arrival time. As mentioned in Section 2.3, iteration-level batching greedily adds arriving requests to the processing batch whenever memory permits. Therefore, the Batching Module tracks memory usage and request arrival times to determine whether new requests can be added to the processing batch. It maintains an active request list that tracks requests that are in the processing batch and the length of their generated tokens. When a request reaches its generation length, it is removed from the batch, freeing memory for new requests. In each iteration, one token is generated per active request, and the active request list is sent to the LLM Serving Simulator (Section 3.4) to estimate the execution time and energy consumption of that iteration. Due to the greedy nature of iteration-level batching, the Batching Module may add requests that exceed memory capacity, as it does not pre-allocate enough memory for the KV cache of generated tokens. In such cases, the most recently added requests and their tokens are temporarily removed to free memory for earlier requests to complete.
3.4. LLM Serving Simulator
The LLM Serving Simulator estimates the execution time and energy consumption for each iteration based on active requests from the Batching Module and Profiling Results. The Profiling Results provide execution times and energy consumptions for key operations such as multi-head attention under various configurations (e.g., number of attention heads, feature dimensions). If a specific data point is missing, the Simulator applies linear interpolation between the nearest profiling data points. We discuss the profiling process in detail in the next subsection and focus here on the Simulator’s execution process. The Simulator estimates the time to process the requests in the active request list. For requests in prefill stage, execution time is estimated by querying the profiling results with the full context length, as the entire sequence is processed in parallel. For requests in decode stage, only the current token is processed, reducing the effective context length to one. If there are $n$ decode requests, execution time is estimated using context length $n$, since all decode requests (each with an effective context length of one) are processed in parallel. The Simulator estimates energy consumption in a similar fashion by querying the profiling results for the energy consumed by each operation. To reduce simulation overhead, the Simulator leverages the repetitive structure of LLMs by simulating only a single Transformer block, which is the smallest set of non-repetitive cells of a LLM (Section 3.2.1). APEX then extrapolates the results to estimate the execution time and energy consumption of the full model. For example, the execution time of the iteration is extrapolated by taking the maximum across all pipeline stages, since the latency is determined by the slowest pipeline stage. In contrast, energy consumption is computed as the sum across all stages, as all devices contribute to energy usage. The Simulator also tracks key LLM serving metrics such as time to first token, time per output token, P95 latency, Model FLOPs Utilization, and Model Bandwidth Utilization; these metrics allow APEX to evaluate the parallel execution plans comprehensively and pick an optimal plan based on a user’s desired metric(s).
3.5. Profiling Results
The profiling results are collected offline and used as system data (Figure 2) during simulation. An Offline Profiler collects platform-specific performance data, which is used by the LLM Serving Simulator to estimate execution time and energy consumption. Instead of profiling each LLM exhaustively, the Profiler performs operation-level profiling, measuring key transformer operations such as multi-head attention and general matrix multiplication. Since these operations are fundamental to transformer models, this approach allows APEX to support various LLMs without requiring per-model profiling. The profiling is performed across various context lengths, attention heads, hidden dimensions, etc. Additionally, the profiler measures collective communication overheads, including AllReduce, ReduceScatter, etc., across different data transfer sizes, device counts, and number of nodes. By covering a wide range of computation and communication configurations, the profiling results facilitate accurate estimation of execution time and energy consumption across different LLMs, device clusters, and parallel execution plans. Notably, profiling is a one-time cost when porting to a new cluster and can be amortized across multiple simulations.
@@ 4. Experimental Results
Summarization Llama-3.1-70B 0.25 1998.82 (1$\times$) 1340.99 (1.49$\times$) 1175.09 (1.70$\times$)
@@
Mistral-Large 0.25 2210.05 (1$\times$) 1624.17 (1.36$\times$) 1514.19 (1.46$\times$)
0.5 1683.00 (1$\times$) 1483.38 (1.14$\times$) 1420.98 (1.18$\times$)
Mixtral-8x22B 0.25 2005.06 (1$\times$) 1753.79 (1.14$\times$) 859.83 (2.33$\times$)
0.5 1618.54 (1$\times$) 1473.79 (1.10$\times$) 856.84 (1.89$\times$)
Creation Llama-3.1-70B 0.25 4027.70 (1$\times$) 2301.54 (1.75$\times$) 1945.75 (2.07$\times$)
0.5 1869.45 (1$\times$) 1639.87 (1.14$\times$) 1133.05 (1.65$\times$)
Mistral-Large 0.25 3561.55 (1$\times$) 2268.50 (1.57$\times$) 1496.45 (2.38$\times$)
0.5 2560.70 (1$\times$) 2016.30 (1.27$\times$) 1313.18 (1.95$\times$)
Mixtral-8x22B 0.25 2645.79 (1$\times$) 2593.91 (1.02$\times$) 1496.45 (1.77$\times$)
0.5 1393.19 (1$\times$) 1386.16 (1.01$\times$) 1035.62 (1.35$\times$)
Chat Llama-3.1-70B 0.25 1622.48 (1$\times$) 1118.85 (1.45$\times$) 824.43 (1.97$\times$)
@@
Mistral-Large 0.25 2041.91 (1$\times$) 1423.44 (1.43$\times$) 1182.75 (1.73$\times$)
0.5 1594.74 (1$\times$) 1356.52 (1.18$\times$) 1102.20 (1.45$\times$)
Mixtral-8x22B 0.25 1938.26 (1$\times$) 1658.22 (1.17$\times$) 575.12 (3.37$\times$)
0.5 1507.11 (1$\times$) 1380.34 (1.09$\times$) 567.53 (2.66$\times$)
4.1. Experimental Setup
To evaluate APEX, we choose a broad range of LLMs and datasets with distinct workloads. We discuss the setup details below.
Models: For evaluation, we choose four state-of-the-art LLMs: Llama-3.1-70B-Instruct, Llama-3.1-405B-Instruct (Dubey et al., 2024), Mistral-Large-Instruct (123B) (Jiang et al., 2023), and Mixtral 8x22B-Instruct (Jiang et al., 2024); This covers a myriad of LLMs in different sizes and also a Mixture-of-Expert (MoE) model. While the models use half-precision (FP16) by default, we also test cases where the model is quantized to FP8 format.
Datasets: We prepared three sets of request traces derived from distinct datasets: a paper summarization dataset (Cohan et al., 2018), a news abstract summarization dataset (Fabbri et al., 2019), and a conversational dataset, LMSYS-Chat-1M (Zheng et al., 2024a). These datasets represent three distinct workloads. The paper summarization (Cohan et al., 2018) exemplifies a prefill-intensive workload, which has a long context length and short generation length. To create a generation-intensive workload with short context and long generation lengths, we modify the news abstract summarization dataset (Fabbri et al., 2019) by appending the following prompt: “Generate a long story using the provided abstract.” This prompts the LLM to produce a long story from a short abstract, resulting in a generation-intensive workload. The LMSYS-Chat-1M (Zheng et al., 2024a) dataset collects real-world conversations between users and LLMs. Most requests in this dataset feature short context and generation lengths, representing a lightweight conversational workload. We randomly subsampled between 512 and 1188 requests from each dataset to produce our request traces. We refer to these traces as Summarization (paper summarization), Creation (news generation), and Chat (LMSYS-Chat-1M) in the following sections. We assume a Poisson distribution (Consul and Jain, 1973) for request arrival times. Details of the request traces are listed in Table 1.
Hardware Platform and Serving Systems: We run APEX simulation using an Intel Xeon 6530 CPU and validated the simulation results against the actual performance of LLM serving on a GPU cluster. We deployed vLLM (Kwon et al., 2023) v0.6.0 and SGLang v0.4.5 (Zheng et al., 2024b) to configure LLM online servers on three GPU clusters: (1) a Single-node H100 cluster equipped with 8 Nvidia H100 SXM GPUs, each with 80 GB of GPU memory, (2) a Multi-node H100 Cluster consisting of two nodes, each with 8 Nvidia H100 SXM GPUs, and (3) a H200 Cluster (single-node) that has 8 Nvidia H200 SXM GPUs.
4.2. Evaluation of APEX’s prediction